OpenStudy (anonymous):

The fraction 3/ (√3a²b) is equivalent to what?

OpenStudy (dumbcow):

\[\frac{3}{\sqrt{3a^{2}b}}=\frac{3\sqrt{3b}}{3ab} = \frac{\sqrt{3b}}{ab}\]